OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 A) x = 0 B) x = 2 C) There is no x value, as there is no solution. D) x can be any value, as there is an infinite number of solutions.

OpenStudy (anonymous):

\(2x+y=1\implies y=-2x+1\) so plug that in to the second equation $$4x+2y=-2\\4x+2(-2x+1)=-2\\4x-4x+2=-2\\\color{red}{2\ne-2}$$oops -- we derived a falsehood by assuming there was a solution, so we can conclude there is no such solution \((x,y)\) satisfying both equations

OpenStudy (anonymous):

@oldrin.bataku so A or C?

OpenStudy (anonymous):

well clearly not A as that suggests there is a solution. there are no solutions